OpenStudy (anonymous):

I have 66% in a class and the final is worth 20%, what do I need to get on the final to get a C (70%) in the class?

OpenStudy (anonymous):

well this is a weighted average. Your current grade is worth 80%. so you want 66 * 0.8 + X * 0.20 = 70

OpenStudy (anonymous):

What percentage grade would I need?

OpenStudy (anonymous):

solve for X

OpenStudy (anonymous):

I'm getting some awkward number.

OpenStudy (anonymous):

show steps and i'll find the problem

OpenStudy (anonymous):

(66 * 0.8)+(70 * .20)=70

OpenStudy (anonymous):

That's link is very complicated Bob06!

OpenStudy (anonymous):

I just need a quick answer.

OpenStudy (anonymous):

66 * 0.8 + X * 0.20 = 70

OpenStudy (anonymous):

solve for X. You put 70 in as X, but X is what you're solving for.

OpenStudy (anonymous):

So I would need an 86 on the final to get a C?

OpenStudy (anonymous):

something like that. just get as good a grade as you can ;) sorry about the complicated link :D

OpenStudy (anonymous):

yep, i get 86

OpenStudy (anonymous):

Thanks!